OpenStudy (anonymous):

find the volume of a regular hexahedron if one of the diagonals of its faces is 8 root 2 inches.

OpenStudy (yacoub1993):

Hint: a "regular hexahedron" is a cube.

OpenStudy (yacoub1993):

hexahedron = volume with six faces. regular = the six faces are the same. The volume of a cube is (side)^3 The diagonal of a face is the hypotenuse of a isosceles right-angle triangle. By Pythagorean theorem: (hypotenuse)^2 = (side)^2 + (side)^2 = 2*(side)^2 (we know all sides of a cube are the same value) (side)^2 = (1/2) * (hypotenuse)^2 (don't forget to square the diagonal before dividing by two) You are given the hypotenuse. Find the value for one side. Then cube that to find the volume.

OpenStudy (yacoub1993):

And Let me know your answer?

OpenStudy (yacoub1993):

@enolreg